A line k passes through the points (4,2) and (-4, 4). A second line h passes through the points (-8,1) and (8, -3). Is line k parallel to line h?
Answer by user_dude2008(1862)

m1=(4-2)/(-4-4)=2/-8=-1/4

m2=(-3-1)/(8--8)=-4/16=-1/4


k is parallel to h